OpenStudy (anonymous):

x + y = 12 x - y = 10 What is the value of the x-determinant for the system shown? a)-22 b)-12 c)-2

OpenStudy (anonymous):

determinant = (1*-1) - (1*1) = -2

OpenStudy (mayankdevnani):

hey!! what are you say about this question..... @omfgitsshelby

OpenStudy (anonymous):

thank you guise C:

OpenStudy (anonymous):

yeah

OpenStudy (mayankdevnani):

plz..tell ,e

OpenStudy (anonymous):

what do you mean?

OpenStudy (mayankdevnani):

plz tell me in simple language!!!

OpenStudy (anonymous):

oh sorry what is the value of x

OpenStudy (anonymous):

x+y=12 x-y=10 2x=22 x=11 11+y=12 y=11